# Monte Carlo Simulation of Operator Dynamics and Entanglement in Dual-Unitary Circuits

### v3

@@ I Introduction
II Locally-scrambled Dual-unitary Dynamics
@@
The unitary $U$ is called dual-unitary when its space-time dual $\tilde{U}$ is also a unitary operator, satisfying $U^{\dagger}U=\tilde{U}^{\dagger}\tilde{U}=I$ [40]. For qubit systems, dual unitary gates are parameterized by four single-qubit unitaries $u_{\pm}$, $v_{\pm}$, and a single parameter $J$:
$U(J)=(u_{+}\otimes u_{-})e^{-i\frac{\pi}{4}(X_{1}X_{2}+Y_{1}Y_{2})-iJZ_{1}Z_{2}}(v_{+}\otimes v_{-}).$ (2)
Here, $\{X_{j},Y_{j},Z_{j}\}$ are the Pauli operators on the $j$-th qubit. Single-qubit gates provide local basis rotations but do not entangle different sites. The non-trivial entanglement arises from the XXZ coupling, where $J\in[0,\pi/4]$ characterizes the scrambling ability of the dual-unitary gate. For $J=\pi/4$, $U$ reduces to a SWAP operation, simply interchanging operators between two sites. Conversely, for $J=0$, it maximally spreading single-site operators to operators supported on both sites.
@@ II.1 Operator Density
$w_{P}(t)~{}\rightarrow~{}w_{P}(t+1)={\sum_{P^{\prime}}}^{\prime}T^{(2)}(P_{x}P_{y},P_{x}^{\prime}P_{y}^{\prime})~{}w_{P^{\prime}}(t),$ (7)
@@
Defining $T(b,b^{\prime})=3^{|b|}T^{(2)}(P,P^{\prime})$, the evolution generated by each two-qubit gate remains
$w_{b}(t)~{}\rightarrow~{}w_{b}(t+1)={\sum_{b^{\prime}}}^{\prime}T(b_{x}b_{y};b_{x}^{\prime}b_{y}^{\prime})~{}w_{b^{\prime}}(t),$ (10)
where the summation $\sum^{\prime}$ is again restricted to the subspace where $b^{\prime}_{z}=b_{z}$ for any $z\neq x,y$. The full Markovian dynamics is then represented by a brick-wall circuit of local transfer matrices $T$, as illustrated in Fig. 1 (a). For the locally-scrambled dual-unitary ensemble, the explicit expression of $T(b,b^{\prime})$ for each two-qubit gate has been computed [41], which reads
$T(b,b^{\prime})=\begin{pmatrix}1&0&0&0\\ 0&0&1-\alpha&\alpha/3\\ 0&1-\alpha&0&\alpha/3\\ 0&\alpha&\alpha&1-2\alpha/3\end{pmatrix}_{bb^{\prime}},$ (11)
with parametrization $\alpha=\frac{2}{3}\cos^{2}(2J)\in[0,2/3]$ that controls the information scrambling power of the dual-unitary gate ensemble: $\alpha=0$ corresponds to the SWAP gate with no scrambling, and $\alpha=1$ corresponds to the fastest scrambling dual unitary gates.
II.2 Operator Entanglement
@@
Under this approximation, we can express the second Rényi entropy in terms of $w_{b}(t)$. As explained in Sec. II.1, the disorder-averaged operator density matrix contains only the diagonal components: $\bar{\rho}=\overline{|O(t)\rangle\langle O(t)|}=\sum_{P}w_{P}(t)|P\rangle\langle P|.$ Tracing out the subregion $\bar{A}$, we obtain the reduced density matrix
$\overline{\rho_{A}}=\sum_{P_{A},P_{\bar{A}}}w_{P_{A}P_{\bar{A}}}(t)~{}|P_{A}\rangle\langle P_{A}|\equiv\sum_{P_{A}}w_{P_{A}}(t)~{}|P_{A}\rangle\langle P_{A}|,$ (14)
where we have decomposed $P=P_{A}\otimes P_{\bar{A}}$ and introduced the marginal distribution $w_{P_{A}}(t)=\sum_{P_{\bar{A}}}w_{P_{A}P_{\bar{A}}}(t)$. The purity of the system is then computed by
$e^{-(n-1)\overline{S^{(n)}_{A}}(t)}\approx\sum_{P_{A}}w^{n}_{P_{A}}(t)=\sum_{b_{A}}\frac{1}{3^{|b_{A}|(n-1)}}w^{n}_{b_{A}}(t),$ (15)
where $b_{A}$ denotes a bit string defined in region $A$.
II.3 Monte Carlo Simulation
Since $w_{b}(t)$ contains an exponentially large number of components, indexed by the bit string $b=\{b_{x}\}$ (with $b_{x}\in\{0,1\}$) that labels the operator support. Direct analytical or numerical study of the evolution of $w_{b}(t)$ following Eq. (10) is impractical. To this end, we would like to use the Monte Carlo method to carry out important sampling and reduce the exponential computation to polynomial complexity. We may interpret the operator support $b$ as an occupation configuration of some fictitious “particles”, and visualize the Markovian dynamics of operator evolution as the evolution of classical particles, where each lattice site $x$ is either empty ($b_{x}=0$) or occupied by a particle ($b_{x}=1$). The system is initialized according to the support of the initial operator $O$. The application of a transfer matrix in Eq. (11) on sites $x$ and $x+1$ is then implemented by the following local update rule, parametrized by the scrambling parameter $\alpha$ of the dual unitary gate:
@@ II.3 Monte Carlo Simulation
as an estimator of Eq. (15). Here, $\delta_{b^{(1)}_{x}b^{(2)}_{x}}$ is the Kronecker delta symbol that enforces $b^{(1)}$ and $b^{(2)}$ have identical configurations inside the entanglement region $A$ of interest, otherwise the quantity $I(b^{(1)},b^{(2)})$ will be zero. Averaging $I(b^{(1)},b^{(2)})$ over many pairs of samples gives $\langle e^{-\overline{S_{A}^{(2)}}}\rangle$, and then $\overline{S_{A}^{(2)}}=-\ln\langle I(b^{(1)},b^{(2)})\rangle$.
@@ III Results
III.1 Operator Density Evolution
@@
Figure 3: Comparing density profile obtained from the recurrence relation in Eq. (20) with Monte Carlo data at $\alpha=0.2$. Initially, panel (a) has two sites occupied in the middle and panel (b) has a single site occupied in the middle. In both panels, solid lines represent the Monte Carlo data with $\sim 10^{7}$ samplings such that the error bar is negligible in the figure. The dashed lines are generated from the recurrence relation Eq. (20). The MF prediction matches well with the MC data with the two-site initial condition at any time $t$, while it deviates from the MC result at large $t$ with a single-site initial condition.
III.1.1 Balanced operator growth from two-site operator
@@
This mean-field approximation neglects any correlation between different sites. The evolution of $\rho(x,t)$ is then derived by applying a layer of transfer matrix $T$ and computing the post-evolution density $\rho(x,t+1)$ [41]. For a transfer matrix $T$ applied to sites $x$ and $x+1$, the evolution reads
$\rho(x,t+1)=\rho(x+1,t)+\alpha\rho(x,t)\Big{(}1-\frac{4}{3}\rho(x+1,t)\Big{)},$ (20)
$\rho(x+1,t+1)=\rho(x,t)+\alpha\rho(x+1,t)\Big{(}1-\frac{4}{3}\rho(x,t)\Big{)}.$
The evolution generated by these equations is plotted in Fig. 3 (a) using dashed lines. The results match the Monte Carlo simulation with high accuracy for $O=Z_{0}Z_{1}$. Additionally, we provide analytical solutions in two limits:
@@
III.1.2 Skewed operator growth from single-site operator
@@
Since $p(t_{0})$ exponentially localizes at small $t_{0}$, evident from Eq. (26), it does not change the long-time relaxation to $\rho(x)=3/4$ and the steady-state operator density near the right light-cone, which is still given by Eq. (25). For the steady-state near the left light-cone, the dependence on $t_{0}$ becomes significant. Defining $\tilde{d}=x+t-2\sim O(1)$, we have $\tilde{\rho}^{\text{LC}}_{0}=\alpha$,
$\tilde{\rho}^{\text{LC}}_{2k}$ $=\frac{3}{4}\sum_{t_{0}=1}^{k}p(t_{0})+p(k+1)$ (29)
$=\frac{3}{4}\left(1-(1-\alpha)^{k-1}\right)+\alpha(1-\alpha)^{k}.$
and
$\tilde{\rho}^{\text{LC}}_{2k+1}$ $=\sum_{t_{0}=1}^{k+1}p(t_{0})\left(\frac{3}{4}-\left(\frac{3}{4}-\alpha\right)(1-\alpha)^{k+1-t_{0}}\right)$ (30)
$=\frac{3}{4}-\frac{3}{4}(1-\alpha)^{k}\left[1-\alpha\left(\frac{4}{3}\alpha+k\left(-1+\frac{4}{3}\alpha\right)\right)\right].$
for any $k>0$.
Again, we test the above prediction near the left light-cone with $O=Z_{1}$ initial condition by directly comparing with Monte Carlo simulations, as exemplified in Fig. 6 (a) and (b) for $k=1,2,3$ at $\alpha=0.2,0.3,0.4$. The operator density near the left light-cone for both left and right movers converges to the mean-field results in Eqs. (29) and (30) as $t$ increases.
Figure 6: Converge of the near-left-light-cone operator density $\tilde{\rho}^{\mathrm{LC}}$ obtained from Monte Carlo sampling with $O=Z_{1}$ initial condition. Panels (a) and (b) display the density evolution of left movers ($\tilde{\rho}_{2k}^{\mathrm{LC}}$) and right movers ($\tilde{\rho}_{2k+1}^{\mathrm{LC}}$) respectively. Colored lines represent the Monte Carlo data with $\sim 10^{7}$ samplings, whereas the horizontal black lines in panels (a) and (b) denote the mean-field results in Eqs. (29) and (30) at corresponding $k$ values, respectively.
III.1.3 Summary of results on operator density
@@
$\tilde{\rho}^{\mathrm{LC}}_{2k}$ Eq. (29) Fig. 6(a)
$\tilde{\rho}^{\mathrm{LC}}_{2k+1}$ Eq. (30) Fig. 6(b)
@@
III.2 Operator Entanglement Growth
Next, we study the dynamics of operator entanglement, focusing on the steady-state entanglement for subregions near the light-cone starting from the single-site operator $O=Z_{1}$. As established in Sec. III.1, the results obtained near the right light-cone for $O=Z_{1}$ should also apply to both light-cone edges for the two-site operator $O=Z_{0}Z_{1}$, so we will not further discuss the operator entanglement dynamics for $O=Z_{0}Z_{1}$, and we will focus our discussion on the case of $O=Z_{1}$ only. Given the initial operator $O=Z_{1}$, as it evolves under the dual unitary circuit, the operator spreads into a light-cone in the spacetime.
@@ III.2 Operator Entanglement Growth
III.2.1 Case 1: from the right light-cone edge
We begin with the Case 1, which allows for a direct mean-field description. With the mean-field approximation Eq. (19), we can estimate the $n$th Rényi operator entanglement Eq. (15) as
$\overline{S^{(n)}_{A,1}}(t)\approx-\frac{1}{n-1}\sum_{x\in A}\ln\left((1-\rho(x,t))^{n}+\rho^{n}(x,t)/3^{n-1}\right).$ (31)
@@ III.2.1 Case 1: from the right light-cone edge
we can safely approximate $\rho^{\text{LC}}_{d}=3/4$. This leads to the expression $\overline{S^{(n)}_{A,1}}\approx(2\ln 2)~{}l_{A}+O(1)$, yielding
@@
III.2.2 Case 2: from the left light-cone edge
@@
Now, let us consider Case 2 near the left light-cone. Similar to the calculation of the operator density, we introduce the conditional joint distribution function $w_{P_{A}}(t|t_{0})$ given an emission of a left-mover for the first time at $t_{0}$ and express the reduced density matrix as follows:
$\overline{\rho_{A}}=\sum_{t_{0}}p(t_{0})\sum_{P_{A}^{(t_{0})}}w_{P_{A}^{(t_{0})}}(t|t_{0})~{}|P_{A}^{(t_{0})}\rangle\langle P_{A}^{(t_{0})}|.$ (33)
@@
Here, $w_{P_{A}}(t|t_{0})$ matches the joint distribution with the initial operator $O=Z_{0}Z_{1}$, with shifted space-time indices. An important observation is that $t_{0}$ determines the position of the leftmost particle. Therefore, for $t_{0},t_{0}^{\prime}\leq(l_{A}+1)/2$ and $t_{0}\neq t_{0}^{\prime}$, we have $\langle P_{A}^{(t_{0})}|P_{A}^{(t_{0}^{\prime})}\rangle=0$. On the other hand, for $t_{0}>(l_{A}+1)/2$, the subsystem $A$ contains only the identity operator $|I\rangle$. Consequently, we can rewrite the reduced density matrix as
$\overline{\rho_{A}}=$ $\sum_{t_{0}\leq\frac{l_{A}+1}{2}}p(t_{0})\sum_{P_{A}^{(t_{0})}}w_{P_{A}^{(t_{0})}}(t|t_{0})~{}|P_{A}^{(t_{0})}\rangle\langle P_{A}^{(t_{0})}|$ (34)
$+\sum_{t_{0}>\frac{l_{A}+1}{2}}p(t_{0})|I\rangle.$
Here, all different terms are orthogonal. Taking the mean-field approximation for the conditional distribution, we can express the operator entanglement entropy as
$e^{-(n-1)\overline{S^{(n)}_{A,2}}}=$ $\sum_{1\leq t_{0}\leq\frac{l_{A}+1}{2}}p^{n}(t_{0})e^{-(n-1)\overline{S^{(n)}_{\tilde{A},1}}}$ (35)
$+\Bigg{(}\sum_{t_{0}>\frac{l_{A}+1}{2}}p(t_{0})\Bigg{)}^{n}.$
In this equation, we have identified the contribution from $\sum_{P_{A}^{(t_{0})}}w_{P_{A}^{(t_{0})}}(t|t_{0})~{}|P_{A}^{(t_{0})}\rangle\langle P_{A}^{(t_{0})}|$ as the operator entanglement of a subsystem $\tilde{A}$, which contains $l_{A}-2t_{0}+2$ sites near the left light-cone for the initial operator $O=Z_{0}Z_{1}$. The corresponding result matches the findings of Case 1 for $O=Z_{1}$.
To extract the volume-law coefficient, we use $\overline{S^{(n)}_{\tilde{A},1}}\approx(l_{A}-2t_{0}+2)\times 2\ln 2$. Moreover, for large $l_{A}$, we can neglect the last term in Eq. (35). The result reads
$e^{-(n-1)\overline{S^{(n)}_{A,2}}}=\alpha^{n}\sum_{1\leq t_{0}\leq\frac{l_{A}+1}{2}}(1-\alpha)^{n(t_{0}-1)}4^{-(n-1)(l_{A}-2t_{0}+2)}.$ (36)
Instead of a direct calculation of the summation, we can estimate by extracting the $t_{0}$ dependence, which scales as $\beta^{t_{0}}$ with $\beta:=2^{4(n-1)}(1-\alpha)^{n}$. When we have $\beta>1$, the dominant contribution is from $t_{0}\approx(l_{A}+1)/2$, while for $\beta<1$, the dominant contribution comes from $t_{0}\approx 0$. Therefore, we have
[…]
@@
This suggests a possible transition in operator entanglement as we tune $\alpha$. Unfortunately, since the mapping from the original quantum circuits to the Markovian process requires $\alpha\leq 2/3$, the transition point $\alpha_{c}=1-2^{-\frac{4(n-1)}{n}}$ is never accessible for integers $n\geq 2$. On the other hand, taking the limit $n=1$ leads to $\alpha_{c}=0$. A non-trivial entanglement transition at finite $\alpha$ only exist for $1<n<(1-(\log_{2}3)/4)^{-1}\approx 1.66$.
@@
III.2.3 Case 3: from the left light-cone edge, focusing on left-movers only

### v4

# myequation

@@ 1 Introduction
2 Locally-scrambled Dual-unitary Dynamics
@@
The unitary $U$ is called dual-unitary when its space-time dual $\tilde{U}$ is also a unitary operator, satisfying $U^{\dagger}U=\tilde{U}^{\dagger}\tilde{U}=I$ [40]. For qubit systems, dual unitary gates are parameterized by four single-qubit unitaries $u_{\pm}$, $v_{\pm}$, and a single parameter $J$: {myequation} U(J)=(u_+⊗u_-)e^-iπ4(X_1X_2+Y_1Y_2)-iJZ_1Z_2(v_+⊗v_-). Here, $\{X_{j},Y_{j},Z_{j}\}$ are the Pauli operators on the $j$-th qubit. Single-qubit gates provide local basis rotations but do not entangle different sites. The non-trivial entanglement arises from the XXZ coupling, where $J\in[0,\pi/4]$ characterizes the scrambling ability of the dual-unitary gate. For $J=\pi/4$, $U$ reduces to a SWAP operation, simply interchanging operators between two sites. Conversely, for $J=0$, it maximally spreading single-site operators to operators supported on both sites.
@@ 2.1 Operator Density
$w_{P}(t)\leavevmode\nobreak\ \rightarrow\leavevmode\nobreak\ w_{P}(t+1)={\sum_{P^{\prime}}}^{\prime}T^{(2)}(P_{x}P_{y},P_{x}^{\prime}P_{y}^{\prime})\leavevmode\nobreak\ w_{P^{\prime}}(t),$ (7)
@@
Defining $T(b,b^{\prime})=3^{|b|}T^{(2)}(P,P^{\prime})$, the evolution generated by each two-qubit gate remains {myequation} w_b(t)  →  w_b(t+1)=∑_b’^′T(b_xb_y;b_x’b_y’) w_b’(t), where the summation $\sum^{\prime}$ is again restricted to the subspace where $b^{\prime}_{z}=b_{z}$ for any $z\neq x,y$. The full Markovian dynamics is then represented by a brick-wall circuit of local transfer matrices $T$, as illustrated in Fig. 1 (a). For the locally-scrambled dual-unitary ensemble, the explicit expression of $T(b,b^{\prime})$ for each two-qubit gate has been computed [41], which reads {myequation} T(b,b’)=(1000001-αα/301-α0α/30αα1-2α/3 )_bb’, with parametrization $\alpha=\frac{2}{3}\cos^{2}(2J)\in[0,2/3]$ that controls the information scrambling power of the dual-unitary gate ensemble: $\alpha=0$ corresponds to the SWAP gate with no scrambling, and $\alpha=1$ corresponds to the fastest scrambling dual unitary gates.
2.2 Operator Entanglement
@@
Under this approximation, we can express the second Rényi entropy in terms of $w_{b}(t)$. As explained in Sec. 2.1, the disorder-averaged operator density matrix contains only the diagonal components: $\bar{\rho}=\overline{|O(t)\rangle\langle O(t)|}=\sum_{P}w_{P}(t)|P\rangle\langle P|.$ Tracing out the subregion $\bar{A}$, we obtain the reduced density matrix
$\overline{\rho_{A}}$ $=\sum_{P_{A},P_{\bar{A}}}w_{P_{A}P_{\bar{A}}}(t)\leavevmode\nobreak\ |P_{A}\rangle\langle P_{A}|$ (12)
$\equiv\sum_{P_{A}}w_{P_{A}}(t)\leavevmode\nobreak\ |P_{A}\rangle\langle P_{A}|,$
where we have decomposed $P=P_{A}\otimes P_{\bar{A}}$ and introduced the marginal distribution $w_{P_{A}}(t)=\sum_{P_{\bar{A}}}w_{P_{A}P_{\bar{A}}}(t)$. The purity of the system is then computed by {myequation} e-(n-1)¯S(n)A(t)≈∑PAwnPA(t)= ∑bA13|bA|(n-1)wnbA(t), where $b_{A}$ denotes a bit string defined in region $A$.
2.3 Monte Carlo Simulation
Since $w_{b}(t)$ contains an exponentially large number of components, indexed by the bit string $b=\{b_{x}\}$ (with $b_{x}\in\{0,1\}$) that labels the operator support. Direct analytical or numerical study of the evolution of $w_{b}(t)$ following Eq. (2.1) is impractical. To this end, we would like to use the Monte Carlo method to carry out important sampling and reduce the exponential computation to polynomial complexity. We may interpret the operator support $b$ as an occupation configuration of some fictitious “particles”, and visualize the Markovian dynamics of operator evolution as the evolution of classical particles, where each lattice site $x$ is either empty ($b_{x}=0$) or occupied by a particle ($b_{x}=1$). The system is initialized according to the support of the initial operator $O$. The application of a transfer matrix in Eq. (9) on sites $x$ and $x+1$ is then implemented by the following local update rule, parametrized by the scrambling parameter $\alpha$ of the dual unitary gate:
@@ 2.3 Monte Carlo Simulation
as an estimator of Eq. (2.2). Here, $\delta_{b^{(1)}_{x}b^{(2)}_{x}}$ is the Kronecker delta symbol that enforces $b^{(1)}$ and $b^{(2)}$ have identical configurations inside the entanglement region $A$ of interest, otherwise the quantity $I(b^{(1)},b^{(2)})$ will be zero. Averaging $I(b^{(1)},b^{(2)})$ over many pairs of samples gives $\langle e^{-\overline{S_{A}^{(2)}}}\rangle$, and then $\overline{S_{A}^{(2)}}=-\ln\langle I(b^{(1)},b^{(2)})\rangle$.
@@ 3 Results
3.1 Operator Density Evolution
@@
Figure 3: Comparing density profile obtained from the recurrence relation in Eq. (3.1.1) with Monte Carlo data at $\alpha=0.2$. Initially, panel (a) has two sites occupied in the middle and panel (b) has a single site occupied in the middle. In both panels, solid lines represent the Monte Carlo data with $\sim 10^{7}$ samplings such that the error bar is negligible in the figure. The dashed lines are generated from the recurrence relation Eq. (3.1.1). The MF prediction matches well with the MC data with the two-site initial condition at any time $t$, while it deviates from the MC result at large $t$ with a single-site initial condition.
3.1.1 Balanced operator growth from two-site operator
@@
This mean-field approximation neglects any correlation between different sites. The evolution of $\rho(x,t)$ is then derived by applying a layer of transfer matrix $T$ and computing the post-evolution density $\rho(x,t+1)$ [41]. For a transfer matrix $T$ applied to sites $x$ and $x+1$, the evolution reads {myequation} ρ(x,t+1)=ρ(x+1,t)+αρ(x,t)(1-43ρ(x+1,t)),ρ(x+1,t+1)=ρ(x,t)+αρ(x+1,t)(1-43ρ(x,t)). The evolution generated by these equations is plotted in Fig. 3 (a) using dashed lines. The results match the Monte Carlo simulation with high accuracy for $O=Z_{0}Z_{1}$. Additionally, we provide analytical solutions in two limits:
@@
3.1.2 Skewed operator growth from single-site operator
@@
Since $p(t_{0})$ exponentially localizes at small $t_{0}$, evident from Eq. (22), it does not change the long-time relaxation to $\rho(x)=3/4$ and the steady-state operator density near the right light-cone, which is still given by Eq. (21). For the steady-state near the left light-cone, the dependence on $t_{0}$ becomes significant. Defining $\tilde{d}=x+t-2\sim O(1)$, we have $\tilde{\rho}^{\text{LC}}_{0}=\alpha$, {myequation} ~ρLC2k=34∑t0=1kp(t0)+p(k+1)=34(1-(1-α)k-1)+α(1-α)k. and {myequation} ~ρLC2k+1=∑t0=1k+1p(t0)(34-(34-α)(1-α)k+1-t0)=34-34(1-α)k[1-α(43α+k(-1+43α))]. for any $k>0$.
Again, we test the above prediction near the left light-cone with $O=Z_{1}$ initial condition by directly comparing with Monte Carlo simulations, as exemplified in Fig. 6 (a) and (b) for $k=1,2,3$ at $\alpha=0.2,0.3,0.4$. The operator density near the left light-cone for both left and right movers converges to the mean-field results in Eqs. (3.1.2) and (3.1.2) as $t$ increases.
Figure 6: Converge of the near-left-light-cone operator density $\tilde{\rho}^{\mathrm{LC}}$ obtained from Monte Carlo sampling with $O=Z_{1}$ initial condition. Panels (a) and (b) display the density evolution of left movers ($\tilde{\rho}_{2k}^{\mathrm{LC}}$) and right movers ($\tilde{\rho}_{2k+1}^{\mathrm{LC}}$) respectively. Colored lines represent the Monte Carlo data with $\sim 10^{7}$ samplings, whereas the horizontal black lines in panels (a) and (b) denote the mean-field results in Eqs. (3.1.2) and (3.1.2) at corresponding $k$ values, respectively.
3.1.3 Summary of results on operator density
@@
$\tilde{\rho}^{\mathrm{LC}}_{2k}$ Eq. (3.1.2) Fig. 6(a)
$\tilde{\rho}^{\mathrm{LC}}_{2k+1}$ Eq. (3.1.2) Fig. 6(b)
@@
3.2 Operator Entanglement Growth
Next, we study the dynamics of operator entanglement, focusing on the steady-state entanglement for subregions near the light-cone starting from the single-site operator $O=Z_{1}$. As established in Sec. 3.1, the results obtained near the right light-cone for $O=Z_{1}$ should also apply to both light-cone edges for the two-site operator $O=Z_{0}Z_{1}$, so we will not further discuss the operator entanglement dynamics for $O=Z_{0}Z_{1}$, and we will focus our discussion on the case of $O=Z_{1}$ only. Given the initial operator $O=Z_{1}$, as it evolves under the dual unitary circuit, the operator spreads into a light-cone in the spacetime.
@@ 3.2 Operator Entanglement Growth
3.2.1 Case 1: from the right light-cone edge
We begin with the Case 1, which allows for a direct mean-field description. With the mean-field approximation Eq. (16), we can estimate the $n$th Rényi operator entanglement Eq. (2.2) as {myequation} ¯S^(n)_A,1(t)≈-1n-1∑_x∈Aln((1-ρ(x,t))^n+ρ^n(x,t)/3^n-1).
@@ 3.2.1 Case 1: from the right light-cone edge
we can safely approximate $\rho^{\text{LC}}_{d}=3/4$. This leads to the expression $\overline{S^{(n)}_{A,1}}\approx(2\ln 2)\leavevmode\nobreak\ l_{A}+O(1)$, yielding
@@
3.2.2 Case 2: from the left light-cone edge
@@
Now, let us consider Case 2 near the left light-cone. Similar to the calculation of the operator density, we introduce the conditional joint distribution function $w_{P_{A}}(t|t_{0})$ given an emission of a left-mover for the first time at $t_{0}$ and express the reduced density matrix as follows: {myequation} ¯ρ_A=∑_t_0p(t_0)∑_P_A^(t_0)w_P_A^(t_0)(t|t_0) |P_A^(t_0)⟩⟨P_A^(t_0)|. Here, $w_{P_{A}}(t|t_{0})$ matches the joint distribution with the initial operator $O=Z_{0}Z_{1}$, with shifted space-time indices. An important observation is that $t_{0}$ determines the position of the leftmost particle. Therefore, for $t_{0},t_{0}^{\prime}\leq(l_{A}+1)/2$ and $t_{0}\neq t_{0}^{\prime}$, we have $\langle P_{A}^{(t_{0})}|P_{A}^{(t_{0}^{\prime})}\rangle=0$. On the other hand, for $t_{0}>(l_{A}+1)/2$, the subsystem $A$ contains only the identity operator $|I\rangle$. Consequently, we can rewrite the reduced density matrix as {myequation} ¯ρA=∑t0lA+12p(t0)∑PA(t0)wPA(t0)(t|t0) |PA(t0)⟩⟨PA(t0)|+∑t0>lA+12p(t0) |I⟩. Here, all different terms are orthogonal. Taking the mean-field approximation for the conditional distribution, we can express the operator entanglement entropy as {myequation} e-(n-1)¯S(n)A,2=1≤t0lA+12pn(t0) e-(n-1)¯S(n)~A,1+(t0>lA+12p(t0))n. In this equation, we have identified the contribution from $\sum_{P_{A}^{(t_{0})}}w_{P_{A}^{(t_{0})}}(t|t_{0})\leavevmode\nobreak\ |P_{A}^{(t_{0})}\rangle\langle P_{A}^{(t_{0})}|$ as the operator entanglement of a subsystem $\tilde{A}$, which contains $l_{A}-2t_{0}+2$ sites near the left light-cone for the initial operator $O=Z_{0}Z_{1}$. The corresponding result matches the findings of Case 1 for $O=Z_{1}$.
@@
To extract the volume-law coefficient, we use $\overline{S^{(n)}_{\tilde{A},1}}\approx(l_{A}-2t_{0}+2)\times 2\ln 2$. Moreover, for large $l_{A}$, we can neglect the last term in Eq. (3.2.2). The result reads {myequation} e^-(n-1)¯S^(n)_A,2=α^n∑_1≤t_0≤lA+12(1-α)^n(t_0-1)4^-(n-1)(l_A-2t_0+2). Instead of a direct calculation of the summation, we can estimate by extracting the $t_{0}$ dependence, which scales as $\beta^{t_{0}}$ with $\beta:=2^{4(n-1)}(1-\alpha)^{n}$. When we have $\beta>1$, the dominant contribution is from $t_{0}\approx(l_{A}+1)/2$, while for $\beta<1$, the dominant contribution comes from $t_{0}\approx 0$. Therefore, we have {myequation} a_2^(n)= {-n2(n-1)ln(1-α) if $\alpha<1-2^{-\frac{4(n-1)}{n}}$,2ln2if $\alpha>1-2^{-\frac{4(n-1)}{n}}$. This suggests a possible transition in operator entanglement as we tune $\alpha$. Unfortunately, since the mapping from the original quantum circuits to the Markovian process requires $\alpha\leq 2/3$, the transition point $\alpha_{c}=1-2^{-\frac{4(n-1)}{n}}$ is never accessible for integers $n\geq 2$. On the other hand, taking the limit $n=1$ leads to $\alpha_{c}=0$. A non-trivial entanglement transition at finite $\alpha$ only exist for $1<n<(1-(\log_{2}3)/4)^{-1}\approx 1.66$.
@@
3.2.3 Case 3: from the left light-cone edge, focusing on left-movers only